A photon has a frequency of 6.64 × 1016 s−1. What is the wavelength of the photon in nm? For your answer, only type in the numerical value with three significant figures. Do NOT include the unit. 1 m = 109 nm, c = 3.0 ×108 m/s
Answer #1

Frequency = 6.64× 10^16 s-1

lambda= c/f= (3.0×10^8 m/s) / ( 6.64×10^16)

= 4.5×10^-9m= 4.5 nm .

Answer = 4.5

{Please give a like}